OpenStudy (anonymous):

Express the radical in simplest form (squ 32)

OpenStudy (anonymous):

to find the simplest radical form, you try to think of any perfect squares that go into 32. 4, 16 are 2. take the larger one, and divide 32 by it. this is what you leave under the square root. outside of it you take the square root of the number you pulled out. ex. \/32 \/2x16 4\/2

OpenStudy (anonymous):

Just think of it like this: 2*16=32 4*8=32 2*16=rad2 rad4. Square rt of 4=2. Hence, 4 rad2. You can also use 4*8 but its a lil extra work. You would get: rad4 rad8=2 rad8=2rad4 rad8. Because 4 is a perfect square you now have 2*2*rad2=4rad2. Hope you understand this.

OpenStudy (anonymous):

\[4\sqrt{2}\]